OpenStudy (anonymous):

Is the horizontal asymptote of: g(x)=x/(x^2+2) zero or non existant?

myininaya (myininaya):

y=0 since the degree of the numerator less than the degree of the denominator

myininaya (myininaya):

is less*

OpenStudy (anonymous):

Doesnt y=1/2?

OpenStudy (amistre64):

HA = 0

myininaya (myininaya):

yes there is a horizontal asymptote at y=0

myininaya (myininaya):

there is no horizontal aymptote when the degree of the numerator is greater than the degree of the denominator

OpenStudy (amistre64):

\[\frac{x}{x^2+2}*\frac{1/x^2}{1/x^2}=\frac{x/x^2}{x^2/x^2 + 2/x^2} \] \[\frac{1/x}{1 + 2/x^2}\] everything with an x in it goes to 0; and we are left with 0/1 = 0

OpenStudy (anonymous):

Sorry for the confusion

OpenStudy (amistre64):

HA is for x at infinity; not x at 0

OpenStudy (amistre64):

VA cant be crossed; HA can be crossed because its only concerned with the infity parts and doesnt care about the innards

myininaya (myininaya):

when the degree of the numerator equals the degree of the denominator you just take the coefficient of the term that has the largest exponent and put it over the coefficient of the term with the largest exponent in the denominator

myininaya (myininaya):

and i missed the words "in the numerator" before the conjunction "and"

OpenStudy (anonymous):

I'm confused. I thought that horizontal (and verticle) asymptotes were attractors, not if x is oo or zero??

myininaya (myininaya):

i don't understand your question

OpenStudy (anonymous):

I know that the HA can be crossed becuase of the hltest

myininaya (myininaya):

hltest?

OpenStudy (anonymous):

nevermind, my computer is really slow so my messages come really late, but Im confused about where the HA is.

OpenStudy (amistre64):

the HA is at y=0

OpenStudy (amistre64):

it is the x axis

myininaya (myininaya):

it is what f(x) gets closer to as x gets really large (or negative large) and if f(x) doesn't get closer to anything then we say there is no horizonal asymptote

OpenStudy (anonymous):

ok, that makes sense. so the HA is 0, right?

myininaya (myininaya):

lol yes my answer doesn't change y=0 is the horizontal asymptote

OpenStudy (anonymous):

I checked this answer and it's not correct (online hmwk). I think that there is no HA.

OpenStudy (anonymous):

wait, actually that's not true either. their is an HA but I have not idea as to what it is.

myininaya (myininaya):

no, y=0 is the horizontal asymptote maybe you type your question in wrong

OpenStudy (anonymous):

It asks y= , and I typed y=0